### v1

# Integrating Reinforcement Learning and Model Predictive Control with Applications to Microgrids

Abstract
This work proposes an approach that integrates reinforcement learning and model predictive control (MPC) to efficiently solve finite-horizon optimal control problems in mixed-logical dynamical systems. Optimization-based control of such systems with discrete and continuous decision variables entails the online solution of mixed-integer quadratic or linear programs, which suffer from the curse of dimensionality. Our approach aims at mitigating this issue by effectively decoupling the decision on the discrete variables and the decision on the continuous variables. Moreover, to mitigate the combinatorial growth in the number of possible actions due to the prediction horizon, we conceive the definition of decoupled Q-functions to make the learning problem more tractable. The use of reinforcement learning reduces the online optimization problem of the MPC controller from a mixed-integer linear (quadratic) program to a linear (quadratic) program, greatly reducing the computational time. Simulation experiments for a microgrid, based on real-world data, demonstrate that the proposed method significantly reduces the online computation time of the MPC approach and that it generates policies with small optimality gaps and high feasibility rates.
keywords:
reinforcement learning , model predictive control , hybrid systems , mixed-integer programming , power systems , mixed-logical dynamical systems , machine learning
[…]
[1]organization=Delft University of Technology, addressline=Melkweg 2, postcode=2628 CD, city=Delft, country=The Netherlands
1 Introduction
1.1 Motivation
Complex infrastructure systems, such as energy, water, and transportation networks are pervasive in our modern world. Analysis and control design of such systems is very challenging due to their size and intricate behavior. These networks inherently have a combination of discrete and continuous dynamics and/or decision variables. Several modeling approaches [1, 2, 3] have considered hybrid systems to represent these important infrastructure networks. For instance, increasing energy demand and expanding decentralized energy generation by renewable sources have motivated the conception of microgrids, which are local electrical grids that can be connected to a main grid. The operation of such microgrids consists of scheduling both the power flow directions and the amount of power being exchanged between each of the entities of the microgrid. As a result, discrete and continuous decision variables have to be planned considering market conditions to reduce the operation cost.
For the considered problem, model predictive control (MPC) arises as a promising technique due to its ability to handle constrained complex systems with discrete and continuous decision variables [4]. In MPC, the system model is used for optimization of the control actions over a finite prediction horizon. At each time step, the model-based optimization problem is solved, the first element of the control sequence is applied to the system, and the prediction horizon is shifted one time step ahead. Furthermore, MPC has solid theoretical foundations with regard to stability, performance, and safety [5]. However, these advantages often come at the price of intense online computational requirements, limiting the use of MPC due to hardware constraints or execution time limitations. This bottleneck is especially more pronounced in MPC for hybrid systems, where the optimizer has also to consider a sequence of discrete decision variables over the given prediction horizon. In this case, the optimization problem then becomes of mixed-integer nature, which is NP-hard [6].
The most widely used technique to solve mixed-integer problems in modern solvers is branch-and-bound [7]. In the worst case, the solver has to find solutions for the relaxed problems for each possible combination of the integer decision variables. This is not scalable with respect to the number of integer variables because of the combinatorial nature of the problem. Branch-and-bound algorithms mitigate this issue by efficiently pruning branches of the search tree by estimating lower and upper bounds of the objective function and by proposing cutting planes, which reduce the feasible set of the mixed-integer programs by introducing linear inequalities as additional constraints [8]. Moreover, many expert-designed heuristics are employed for improving the search such as node selection and branch variable selection techniques to reduce computation time. One alternative to branch-and-bound for the control of hybrid systems is to pre-compute the MPC control law offline and simply evaluate this function online. For hybrid systems, the off-line computation of the explicit MPC control law via multi-parametric programming and dynamic programming was explored in [6]. However, this approach, also referred to as explicit MPC, can be only successfully applied to low-dimensional linear systems. Despite improvements in branch-and-bound and explicit MPC, solving mixed-integer problems with a significant number of integer decision variables is fundamentally difficult.
Recently, supervised learning has been explored in several MPC approaches for hybrid systems to reduce the online computational time of the resulting mixed-integer programs [9, 10, 11, 12]. In essence, the aforementioned supervised learning methods have the same structure and learning setting and the main difference lies in the choice of the classifier. These works employ supervised learning to approximate the mapping from the system state to the discrete optimal solution. By applying this classifier to predict the discrete optimization variables, the mixed-integer program is then simplified to an optimization problem consisting only of real-valued variables, significantly reducing the online computational burden of the MPC controller. To build the training dataset, these methods rely on branch-and-bound to solve the control problem – a mixed-integer program – to optimality several times. As a result, the main issue is sidelined to the offline phase of the algorithm, where typically more computing resources are available. Moreover, in supervised learning the goal is to reduce the classification error, e.g. the distance between the predicted discrete sequence and the optimal discrete sequence. Even though the classification error is typically a suitable proxy for control performance, this might not necessarily be true depending on the mapping from the input sequence to the control objective function. For a more complete overview of the literature on the intersection of learning and control for hybrid systems, the reader is referred to Section 2.
1.2 Contributions
Our work aims to propose an integrated reinforcement learning and MPC method that solves mixed-integer linear and quadratic programs with low computational footprint, low optimality gap, and high feasibility rate. We build on the existing idea of decoupling the decision on the discrete and continuous variables with learning and MPC [9, 10, 11, 12]. However, we explore a novel paradigm by employing reinforcement learning in place of supervised learning to directly optimize for control performance – instead of minimizing the classification error – and to avoid the use of branch-and-bound in both the offline and online phases of the algorithm.
The main contributions of the paper with regard to the literature are:
  • 1.
    We propose a novel integrated reinforcement learning and MPC framework for control of mixed-logical dynamical systems.
  • 2.
    The Q-function is partitioned across the prediction horizon and the definition of decoupled Q-functions is conceived to make the learning problem more tractable.
  • 3.
    Simulation experiments in a microgrid system show the efficacy of the proposed approach in reducing the computational load of the MPC controller. Moreover, the comparison between the proposed approach and a method based on supervised learning reveals a trade-off: while the former outperforms in terms of feasibility, the latter outperforms in terms of optimality.
In our problem, long-term decision-making involves determining suitable sequences of discrete decision variables over a given prediction horizon, which is a combinatorial problem. To mitigate the combinatorial complexity of the optimization problem, we propose the decoupling of the Q-function across time, i.e., each time step of the prediction horizon has its own decoupled Q-function. With the proposed decoupling of the Q-function, our method effectively simplifies the learning problem as the discrete actions are decided sequentially rather than simultaneously. A similar idea was explored in [13]; however, the separation therein was not in time, but for independent actions. In our setting, future actions may depend on past actions; therefore, we use a recurrent neural network to retain past information in its internal states and recursively compute the decoupled Q-functions. A case study is used to empirically show that the reinforcement learning agent can learn performing policies with the unorthodox definition of decoupled Q-functions.
1.3 Outline
This paper is organized as follows. In Section 2, we give an overview of the literature on the intersection of learning and control for hybrid systems. Section 3 formalizes the control problem. Our novel method that integrates reinforcement learning and MPC for the control of mixed-logical dynamical systems is described in Section 4. Section 5 presents the simulation setup, the results, and the discussion. This paper ends in Section 6 with conclusions and suggestions for future work.
2 Related work
Some works have applied learning to reduce the solution time of mixed-integer programs by embedding learning into the branch-and-bound algorithm to substitute expert-designed heuristics with learned rules. For instance, learning can be used to improve cutting plane rules [14], branching variable selection [15], and node selection [16]. Although the aforementioned approaches indeed reduce the solution time of the considered mixed-integer programs, control applications often require a more expressive reduction in the computational load.
In [17] the authors approximate the state-action value function (Q-function) of time-invariant mixed-logical dynamics systems. The core of their approach is to parametrize the Q-function with Benders cuts and estimate the Q-function from a lower bound. However, the scope of application of this approach is limited because it only considers auxiliary binary variables from the mixed-logical dynamical modeling framework and it is not compatible with binary states and inputs. In [18], the authors are successful in reducing the computational burden of MPC by learning a state-dependent horizon and a state-dependent recomputation policy, i.e. the policy that decides whether the MPC optimal control problem should be recomputed at a given time step. In [19], an MPC controller is used as a function approximator of the state-action value. The authors show how the internal parameters of the controller can be tuned by policy gradient methods. An extension of the same approach to mixed-integer problems was made in [20]; however, the goal is to target performance rather than to reduce the computational cost of the optimization problem.
There is also a body of work that explores the use of reinforcement learning to jointly learn discrete and continuous policies [21, 22, 23]. In the artificial intelligence community, this problem is explored under the framework of Markov decision processes with parameterized actions. The problem is very similar to that of control of hybrid systems because the action space has discrete and continuous elements. Even though these works can provide useful insight into the parametrization and training of policies for hybrid systems, they lack the optimality and constraint satisfaction that MPC can provide.
Outside the domain of hybrid systems, other works have also integrated learning into the MPC framework in various ways. The most popular approaches include the adaptation of the system model, the use of MPC as a safety filter for RL, and the online tuning of the cost and constraint functions for performance, see [24, 25] for more methods on the interplay of learning and MPC. These approaches are aimed towards improving performance, safety, and/or robustness and do not address the computational issues, which is the main concern of our work.
3 Control problem
As a control algorithm, we consider MPC for its capacity to handle multivariable constrained hybrid systems. In the context of MPC, mixed-logical dynamical (MLD) systems [26] are typically used to model hybrid systems since they can be readily used to formulate open-loop finite-horizon optimal control problems [4]. The equivalence of MLD systems and other hybrid system modeling frameworks was established in [27], showing their broad applicability. When MLD systems are controlled with MPC and the cost function and the constraints are linear (quadratic), the resulting optimization problem is a mixed-integer linear (quadratic) program (MIL(Q)P). In this section, a general description of the MPC optimization problem for an MLD system is first addressed to set the stage for the formulation of the control problem as an MIL(Q)P.
Consider the MLD system
$\begin{split}x(k+1)&=Ax(k)+B_{1}u(k)+B_{2}\delta(k)+B_{3}z(k)+B_{5},\\ &E_{2}\delta(k)+E_{3}z(k)\leq E_{1}u(k)+E_{4}x(k)+E_{5}\end{split}$ (1)
where $x\in\mathbb{R}^{n_{\mathrm{c}}}\times\{0,1\}^{n_{\mathrm{d}}}$ is a vector containing the continuous and discrete system states, $u\in\mathbb{R}^{m_{\mathrm{c}}}\times\{0,1\}^{m_{\mathrm{d}}}$ are the continuous and discrete inputs, $\delta\in\{0,1\}^{r_{\mathrm{d}}}$ is a vector with the auxiliary discrete variables, $z\in\mathbb{R}^{r_{\mathrm{c}}}$ is a vector with the continuous auxiliary variables arising from the MLD modeling, and $A,\ \{B_{i}\}_{i=1,2,3,5},\ \{E_{i}\}_{i=1,2,3,4,5}$ are matrices of appropriate dimensions. Note that the linear constraints may represent logical constraints, a byproduct of MLD modeling, or system operating constraints.
The goal is to control the system (1) with a receding-horizon strategy, where at each time step a finite-horizon optimal control problem is solved. Such an optimization problem can be formulated as follows:
$\begin{split}\min_{\mathbf{x}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k)}\ &J(\mathbf{x}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k))\\ \text{s.t. }&x(k+l+1)=Ax(k+l)+B_{1}u(k+l)+\\ &\hskip 28.45274pt+B_{2}\delta(k+l)+B_{3}z(k+l)+B_{5},\\ &E_{2}\delta(k+l)+E_{3}z(k+l)\leq E_{1}u(k+l)+\\ &\hskip 85.35826pt+E_{4}x(k+l)+E_{5},\\ &\text{for }l=0,\dots,N_{\mathrm{p}}-1\end{split}$ (2)
where the cost function is defined by
$\begin{split}&J(\mathbf{x}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k))=\\ =&\sum_{l=0}^{N_{\mathrm{p}}-1}\left(\ell(x(k+l),u(k+l),\delta(k+l),z(k+l))\right)+\\ &\hskip 85.35826pt+V_{\mathrm{f}}(x(k+N_{\mathrm{p}})),\end{split}$ (3)
the prediction horizon is denoted by $N_{\mathrm{p}}$, the variable $l$ indexes the time along the prediction horizon, and the final cost is represented by $V_{\mathrm{f}}(\cdot)$. The predicted state trajectory over the prediction horizon is denoted by $\mathbf{x}(k)=[x^{\mathrm{T}}(k),\ldots,x^{\mathrm{T}}(k+N_{\mathrm{p}})]^{\mathrm{T}}$. Let the predicted input be explicitly divided into continuous and discrete components $u(k)=[u_{\mathrm{c}}^{\mathrm{T}}(k),\ u_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}$. The stacked continuous input and auxiliary variables over the prediction horizon are represented in vector $\epsilon_{\mathrm{c}}(k)=[u_{\mathrm{c}}^{\mathrm{T}}(k),z^{\mathrm{T}}(k),\ldots,u_{\mathrm{c}}^{\mathrm{T}}(k+N_{\mathrm{p}}-1),\ z^{\mathrm{T}}(k+N_{\mathrm{p}}-1)]^{\mathrm{T}}$. Similarly, the stacked discrete inputs and auxiliary variables over the prediction horizon are expressed by $\epsilon_{\mathrm{d}}(k)=[u_{\mathrm{d}}^{\mathrm{T}}(k),\delta^{\mathrm{T}}(k),\ldots,u_{\mathrm{d}}^{\mathrm{T}}(k+N_{\mathrm{p}}-1),\ \delta^{\mathrm{T}}(k+N_{\mathrm{p}}-1)]^{\mathrm{T}}$. For the sake of simplicity, the constraints are assumed to be polyhedral as in (1).
At each time step, the optimization (2) is solved and the first entry of the decision variables is applied to the system. Let $(\mathbf{x}^{*}(k),\ \epsilon_{\mathrm{c}}^{*}(k),\ \epsilon_{\mathrm{d}}^{*}(k))$ be the optimal solution of (2). The MPC control law then is defined by
$u_{\mathrm{MPC}}(k)=[\epsilon_{\mathrm{c},0}^{*,T}(k),\ \epsilon_{\mathrm{d},0}^{*,T}(k)]^{\mathrm{T}}$ (4)
where the first entries of the solution vector are expressed as $[\epsilon_{\mathrm{c},0}^{*,T}(k),\ \epsilon_{\mathrm{d},0}^{*,T}(k)]^{\mathrm{T}}=[u_{\mathrm{c}}^{*,T}(k),\ z^{*,T}(k),\ u_{\mathrm{d}}^{*,T}(k),\ \delta^{*,T}(k)]^{\mathrm{T}}$.
If the cost function (3) is quadratic (linear), then the optimization problem (2) can be conveniently recast as a mixed integer quadratic (linear) program [4]. The corresponding quadratic optimization problem in a compact form can be stated as follows:
$\begin{split}\min_{\epsilon(k)}\ &\epsilon^{\mathrm{T}}(k)H_{1}\epsilon(k)+\epsilon^{\mathrm{T}}(k)H_{2}x(k)+c^{\mathrm{T}}\epsilon(k)\\ \text{subj. to }&G\epsilon(k)\leq w+Sx(k)\end{split}$ (5)
where $\epsilon(k)=[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}$, and $H_{1},\ H_{2},\ c,\ G,\ w,\ S$ are matrices of appropriate sizes. If the cost function is linear, then the matrices $H_{1}$ and $H_{2}$ are zero.
4 Method: Novel integration of reinforcement learning and MPC
In this section, we present the main contribution of our work: an integrated reinforcement learning and MPC method for the solution of mixed-integer linear and quadratic problems for the control of mixed-logical dynamical systems. The main goal is to ease the online computational burden of solving mixed-integer linear and quadratic programs by decoupling the computation of the integer and continuous decision variables. In our framework, the integer decision variables are determined by reinforcement learning, and the continuous decision variables by MPC, which is an optimization-based framework. Accordingly, the solution of mixed-integer programs is entirely avoided in on-line operation. Consequently, our approach can be seen as an alternative to branch-and-bound for MPC for MLD systems. In what follows, the decoupling of the discrete and continuous variables of the MILP of (5) is addressed. Next, we explain the role of reinforcement in our approach.
Herein it is assumed that the system dynamics are known and training takes place fully offline, i.e., there is no online adaptation of the learning parameters.
4.1 Decoupling of decision variables
The mixed-integer linear program (MILP) (5) is parametric with respect to the initial state $x(k)$ and possibly an exogenous signal $\gamma(k)$. Consider the set of these values in which (5) is feasible. In this set, the solution of (5) maps the signal $(x(k),\ \gamma(k))$ to their corresponding optimal decision variables $(\epsilon_{\mathrm{c}}(k)^{*},\ \epsilon_{\mathrm{d}}(k)^{*})$. Note that the uniqueness of the solution of (5) is not assumed and multiple optima may exist. In this case, only one solution is considered. Henceforth, we use the augmented state
$\chi(k)=[x^{\mathrm{T}}(k),\ \gamma^{\mathrm{T}}(k)]^{\mathrm{T}}$
to refer to the augmented state $\chi(k)$ containing the system state $x(k)$ and the exogenous signal $\gamma(k)$ at time step $k$.
Consider the scenario where the discrete part of the optimal solution $\epsilon_{\mathrm{d}}(k)^{*}$ is known. Then the continuous part of the solution $\epsilon_{\mathrm{c}}(k)^{*}$ can be easily recovered by solving the problem (5), which becomes a linear program (LP) with the fixed discrete optimization variables. The principle used in our approach to decouple the discrete and continuous decision variables is based on this fact. In the proposed approach, the discrete variables are determined by reinforcement learning. Then solving for the continuous MILP (5) can be simplified into the following LP:
$\begin{split}V(\chi(k),\ \epsilon_{\mathrm{d}}(k))=\min_{\epsilon_{\mathrm{c}}(k)}\ &c_{\mathrm{c}}^{\mathrm{T}}\epsilon_{\mathrm{c}}(k)\\ \text{subj. to }&G[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}\leq w+S\chi(k)\end{split}$ (6)
The main difficulty of solving MILPs is the discrete optimization variables and modern solvers mostly rely on branch-and-bound as a strategy to address this problem. Therefore, computing these complicating decision variables with learning-based methods removes the main optimization hurdle. Furthermore, solving such LPs of the form (6) is computationally cheap with state-of-the-art specific-purpose solvers. The immediate benefit of the continuous variables being determined by an optimization-based control, such as MPC, is constraint satisfaction and optimality – given the feasibility of (6). A depiction of the decoupling process is shown in Figure 1.
Figure 1: Representation of the decoupling of the discrete and continuous decision variables. From mixed logical dynamical (MLD) modeling and the use of an MPC approach for control, a mixed-integer linear program (MILP) can be formulated for the operation of the microgrid. Then, a learning approach – either reinforcement learning (RL) or supervised learning (SL) – is used to determine the discrete variable $\epsilon_{\mathrm{d}}(k)$. The MILP is then simplified into a linear program (LP), which computes the continuous variable $\epsilon_{\mathrm{c}}(k)$.
4.2 Role of reinforcement learning
Reinforcement learning (RL) is a general learning framework where the agent learns a control policy based on its interaction with the environment. It has received increasing attention in control applications due to its capacity to learn complex policies and for its low demand for online computation [28]. Here we describe an approach that exploits the benefits of RL to efficiently determine the discrete decision variables of (5).
In the mixed-integer linear program (5), the main computational complexity stems from the number of discrete decision variables, which can come from both the number of actions per time step and the length of the prediction horizon. In our method, reinforcement learning is used to ease the online computational burden of solving mixed-integer linear programs by decoupling the computation of the discrete and continuous decision variables. In our framework, the discrete decision variables are determined by reinforcement learning, and the continuous decision variables by optimization-based control – MPC. Accordingly, the solution of mixed-integer programs is entirely avoided in on-line operation.
In order to formulate the problem as a Markov decision process (MDP) – the standard reinforcement learning framework – we lump together the system and the MPC controller in a single block to form the environment, see Figure 2. The environment receives the discrete actions from the RL agent and outputs the next state and the corresponding reward. This abstraction allows the decoupling of the discrete and continuous decision variables and it is crucial to bridge reinforcement learning and optimization-based control for MLD systems. The agent is responsible for determining the discrete sequence of actions $\epsilon_{\mathrm{d}}(k)$ over the prediction horizon. This vector is then sent to the environment which solves the optimization problem (6) and computes the continuous decision variables $\epsilon_{\mathrm{c}}(k)$. As a result, the environment outputs the next state $\chi(k+1)$ and the reward $r(k)$. Any reinforcement learning algorithm can be employed to train the agent; however, for simplicity, we use Deep Q-Learning [29] to present our approach. For a graphical representation of the reinforcement learning setting, see Figure 2.
Figure 2: A depiction of the proposed control scheme that integrates reinforcement learning into an MPC framework. The agent’s goal is to maximize its long-term reward. It learns to adapt its policy by repeatedly interacting with the environment, that is, by sending a discrete action $\epsilon_{\mathrm{d}}(k)$ and by receiving the extended state $\chi$ and immediate reward $r$. The MPC controller, which is lumped in the environment, receives this discrete action $\epsilon_{\mathrm{d}}(k)$ and then solves an optimization problem to determine the continuous action $\epsilon_{\mathrm{c}}(k)$. Finally, the input $\epsilon$ is fed to the system and the next state is computed.
Herein, the Q-function is defined as the expected reward over a finite horizon if the agent takes action $\epsilon_{\mathrm{d}}(k)$ at the current time step $k$ and then follows the policy $\epsilon_{\mathrm{d}}(k+l)=\pi(\chi(k+l))$ over the remaining steps of the horizon:
$Q^{\pi}(\chi(k),\epsilon_{\mathrm{d}}(k))=r(k)+\sum_{l=1}^{N_{\mathrm{p}}-1}\alpha^{l}r(k+l)$ (7)
where $\alpha$ is the discount factor and the reward is defined as a function of the objective of (6):
$r(k)=f(V(\chi(k),\ \epsilon_{\mathrm{d}}(k)))$
[…]
The vector $\epsilon_{\mathrm{d}}(k)$ contains the entire sequence of discrete variables over the prediction horizon. For each time step of the prediction horizon $l=0,\ldots,N_{\mathrm{p}}-1$, we can represent the discrete action per time step with $\varepsilon_{\mathrm{d},l}(k)$, hence
$\epsilon_{\mathrm{d}}(k)=[\varepsilon_{\mathrm{d},0}^{\mathrm{T}}(k),\ \varepsilon_{\mathrm{d},1}^{\mathrm{T}}(k),\ \ldots,\ \varepsilon_{\mathrm{d},N_{\mathrm{p}}-1}^{\mathrm{T}}(k)]^{\mathrm{T}}\ .$
For example, suppose that all of the discrete optimization variables are binary and the parameter $N_{\mathrm{b}}$ represents the number of binary actions per time step, then there are $2^{N_{\mathrm{b}}\cdot N_{\mathrm{p}}}$ possible sequences for the vector $\epsilon_{\mathrm{d}}(k)$. As the number of binary variables $N_{\mathrm{b}}$ and the length of the prediction horizon $N_{\mathrm{p}}$ increase, the number of sequences grows exponentially. Due to the potentially large size of the action space, this is a very challenging problem for RL algorithms. For instance, efficient exploration becomes a big issue in this setting.
Instead of using an RL algorithm to find the Q-function for the action $\epsilon_{\mathrm{d}}(k)$, as defined in (7), our approach shifts the goal to learning the so-called decoupled Q-functions for the sub-actions $\{\varepsilon_{\mathrm{d},l}(k)\}_{l=0}^{N_{\mathrm{p}}-1}$. The decoupled Q-functions are pseudo Q-functions in the sense that they do not entirely fit the standard definition of a Q-function. These decoupled Q-functions are just used as proxies to determine the sub-actions, i.e.,
$\varepsilon_{\mathrm{d},l}(k)=\max_{\varepsilon}\ Q_{l}(\chi(k),\ \varepsilon,\ h_{l-1})\ {\text{for }l=1,\ \ldots,\ N_{\mathrm{p}}-1}.$ (8)
where $h_{l-1}$ is a signal that depends on the previous actions up to time step $i$ of the prediction horizon $\{\varepsilon_{\mathrm{d},j}(k)\}_{j=0}^{j=i}$. If the sub-actions were independent, the learning problem would be greatly simplified. However, this assumption is very constraining for most useful applications where optimal planning requires knowledge of past sub-actions. While each step of the prediction horizon has its own decoupled Q-function $Q_{l}(\cdot)$, the signal $h_{l-1}$ retains some of the couplings with the past sub-actions. In this way, some inter-dependency between the sub-actions is preserved and the learning problem is simplified with the decoupling of the Q-functions. Note that the size of the action space for each decoupled Q-function is $2^{N_{\mathrm{b}}}$, showing that the combinatorial growth is mitigated with respect to the prediction horizon.
To approximate the decoupled Q-functions, we propose the use of recurrent neural networks (RNNs), which have internal states that carry over information from the past – similar to the role of the signal $h$ in (8). More specifically, a long short-term memory (LSTM) network is employed for its capacity to capture long-term time dependencies in data [30]. It is important to stress that the LSTM is a particular implementation of (8) and similar RNNs can also be used to attain the same objective. An unrolled LSTM is shown in Figure 3 to represent the approximation of the decoupled Q-values. The idea of using an LSTM network to decouple the actions of each time step of the prediction horizon was first explored in [11] in the context of supervised learning. Our approach, however, employs the LSTM network as a function approximator of the decoupled Q-functions in the setting of reinforcement learning.
It is worth stressing that the decision-making between time steps is not fully decoupled because the internal state of the LSTM network carries information forward in time, maintaining the dependence of future actions on past actions. Moreover, the integer and continuous decision variables are also not entirely decoupled because the learning agent is trained based on the augmented state and reward signals, which depend on all the decision variables. Therefore, in both cases, decoupling aids tractability while maintaining some degree of interconnection, which is arguably desirable in our setting.
Figure 3: A representation of the recurrent LSTM network on the left-hand side and the unrolled LSTM network on the right-hand side. Note that the LSTM is unrolled for the duration of the prediction horizon $N_{\mathrm{p}}$. Moreover, at each time step $k$, the augmented state $\chi(k)$ can go through the operation $g_{k}(\cdot)$, changing the manner in which the augmented state is presented at the LSTM network at time step $k$. This can be interpreted as a pre-processing technique to better exploit the structure of our problem.
Each episode of the training procedure starts with a randomly chosen initial state $\chi(k)$. Then, the RL agent repeatedly proposes a discrete action $\epsilon_{\mathrm{d}}(k)$, sends it to the environment, and collects the reward $r(k)$ and the next state $\chi(k+1)$. An episode ends when a pre-determined time limit is reached or when the RL agent suggests a discrete action that makes the optimization problem (6) infeasible. At each time step, the transitions $(\chi(k),\ \epsilon_{\mathrm{d}}(k),\ r(k),\ \chi(k+1))$ are stored in a data buffer $\mathcal{D}$, which only holds a finite number of most recent transitions. After the end of each episode, $T$ transitions are randomly sampled from $\mathcal{D}$ and the weights of the LSTM network are updated by gradient descent on the following loss function:
$L=\frac{1}{T}\sum_{k=1}^{T}\sum_{l=0}^{N_{\mathrm{p}}-1}(y(k)-Q_{l}(\chi(k),\ \varepsilon_{\mathrm{d},l}(k)))^{2}$
where the temporal-difference (TD) target is defined as
$y(k)=r(k)+\alpha\frac{1}{N_{\mathrm{p}}}\sum_{l=0}^{N_{\mathrm{p}}-1}\max_{\varepsilon}Q_{l}(\chi(k),\ \varepsilon)$
and the dependency of the decoupled Q-functions on $h$ has been omitted for the sake of simplicity. The summation on the right-hand side represents the average of the maximum decoupled Q-values for each of the sub-actions. Note that a single TD target is used for all decoupled Q-functions. Some alternative definitions of the TD target can be found in [13]. Training ends when a pre-determined number of episodes or number of transitions is reached.
Once training is completed, the RL agent is employed to suggest the discrete action $\epsilon_{\mathrm{d}}(k)$ and, consequently, the MPC optimization problem (2) is simplified to the LP described in (6). As a result, the solution time of the MILP is replaced by the time of inference of an LSTM network and the solution time of an LP, which can be efficiently solved by modern solvers, greatly reducing the computation time.
4.3 Differences between methods based on supervised and reinforcement learning
To shed light on the fundamental differences between methods based on supervised learning and our approach, we explain the main differences in the learning setups. From the perspective of supervised learning approaches, the problem becomes a classification problem, i.e. the network aims to predict the optimal sub-action for each time step of the prediction horizon. Instead of relying on decoupled Q-values to obtain the sub-actions, SL directly approximates the mapping from the augmented state to the optimal sub-actions. Evidently, this comes at the cost of repeatedly solving the MILP (5) to optimality to create the training dataset. Moreover, in SL, training does not aim at minimizing the long-term system’s operation cost as in RL, rather it targets at minimizing the classification error. The difference in the objective of RL and SL creates a notable difference in the behavior of the control policies, which are best shown in the case study of the next section. For a brief description of the training process in the context of supervised learning, the reader is referred to Appendix B.
5 Case study: microgrid control
In this section, we evaluate the performance of our approach for microgrid operation optimization. The goal is to minimize the long-term operation cost, typically defined over several hours, of the microgrid system described in Figure 4 and Appendix A. At each time step, the microgrid operator can determine a set of discrete and continuous decision variables, shown in Table 1. Herein we consider the receding horizon control strategy laid out in Section 3. In this case, the microgrid operator has to determine a sequence of discrete and continuous decision variables over a finite prediction horizon. Then, the underlying optimization problem is formulated as an MILP (15). Hence, efficiently solving such an MILP to minimize the required computation time is at the core of our problem. In this context, we compare our RL-based approach against standard MPC for MLD systems, where the optimization problem (2) is solved to optimality with branch-and-bound, and against the supervised learning approach described in the last section. To keep the comparisons fair, we use the general-purpose solver GUROBI [31] for the corresponding LPs and MILPs. Besides, as performance metrics, we analyze the total operation cost, feasibility rate, and computation time.
Figure 4: Depiction of the elements of a microgrid and a bidirectional connection with the main grid. The pointed arrows indicate the possibility of power flow between two elements.
5.1 Setup
In the case study, we aim to solve the economic power dispatch problem for a microgrid, i.e. to fulfill the power demand of the microgrid loads while minimizing the operation cost. The knowledge of the market energy prices and the forecasts of the power demand and power generation from renewable energy sources – solar and wind energy are considered – allows the operator to plan and optimally schedule the operation of the microgrid over a time horizon. To accomplish this goal, the microgrid operator has a set of discrete and continuous actions. The discrete control actions are the following: buying or selling from grid $\delta_{\mathrm{grid}}(k)$, charging or discharging the energy storage system $\delta_{\mathrm{b}}(k)$ and turning on or turning off the generators $\{\delta_{i}^{\mathrm{dis}}(k)\}_{i=1}^{N_{\mathrm{gen}}}$ at time step $k$. The continuous actions are the power exchange with the main grid $P_{\mathrm{grid}}(k)$, power exchange with the energy storage system $P_{\mathrm{b}}(k)$, and the power provided by the dispatchable generators $P_{i}^{\mathrm{dis}}(k)$. For a summary of the decision variables, see Table 1. In this setting, the decision variables are represented by the vectors $\epsilon_{\mathrm{c}}(k)$ and $\epsilon_{\mathrm{d}}(k)$, which are formed by the concatenation of all discrete and continuous decision variables, respectively, over the prediction horizon. Furthermore, consider the exogenous signal $\gamma(k)$ contains the forecasts over the prediction horizon for the: purchasing, sale, and production prices; power generation from the renewable energy sources; and power demand from the loads.
Variable Description
$\delta_{\mathrm{b}}$ status of the storage unit:
charging(1)/discharging(0)
$\delta_{\mathrm{grid}}$ mode of connection with main grid:
importing(1)/exporting(0)
$\delta_{i}^{\mathrm{dis}}$ state of dispatchable unit $i$: on(1)/off(0)
$P_{\mathrm{b}}$ power exchanged with the storage unit [kW]
$P_{\mathrm{grid}}$ power exchanged with the main grid [kW]
$P_{i}^{\mathrm{dis}}$ power generated by dispatchable unit $i$ [kW]
Table 1: Decision variables and their descriptions.
The parameters of the microgrid considered in this case study are shown in Table 3. For the experiments, the number of dispatchable generators is set to $N_{\mathrm{gen}}=3$. Hence, for a prediction horizon $N_{\mathrm{p}}$, there are $5\cdot N_{\mathrm{p}}$ binary variables. As there is a binary variable for the connection with the main grid and another one for the connection with the storage unit, amounting to $5$ binary variables per time step. Hence, even for a small number of generators, the number of possible action sequences ($2^{5\cdot N_{\mathrm{p}}}$) is potentially huge depending on the chosen prediction horizon. With a sampling time of $T_{\mathrm{s}}=30$ minutes, all the methods are simulated with a prediction horizon of $N_{\mathrm{p}}\in\{4,12,24,48\}$, corresponding to look-ahead periods of 2, 6, 1226122,\ 6,\ 122 , 6 , 12, and $24$ hours, respectively. The objective of these experiments with a varying prediction horizon is to assess how the effectiveness of the methods scales with the increase of the number of binary variables.
The renewable energy generation and load profiles were taken from the actual operation of the Dutch main grid, which is publicly available at [32]. Since the microgrid considered in this case study does not have the capacity to meet the demand of the entire grid, the real power profiles were linearly down-scaled to adjust the values to our case study. This is the only pre-processing operation applied to these signals, so the shapes of the renewable energy generation and load profiles remain unchanged. The RL-based and SL-based algorithms are trained on one year of data (2022) and their performances are evaluated in a different year (2021) to assess the capacity of the approaches to generalize and adapt to a new setting. The price profiles – $c_{\mathrm{prod}},\ c_{\mathrm{buy}},\ c_{\mathrm{sell}}$ – are not publicly available and, consequently, they were synthetically generated by three different normal distributions which obey a basic principle: on average producing energy is cheaper than purchasing energy from the main grid and higher than selling energy to the main grid. Due to the randomness of the prices, which mimic real market conditions, the price order can change and the microgrid operator can take advantage of some occasions. For instance, when the production price is smaller than the sell price, it is profitable to generate power that exceeds the load demand and sell the power surplus to the main grid.
Prediction horizon $N_{\mathrm{p}}=4$ $N_{\mathrm{p}}=12$ $N_{\mathrm{p}}=24$ $N_{\mathrm{p}}=48$
Learning Approach SL RL SL RL SL RL SL RL
Optimality gap 0.08% 0.79% 0.13% 1.73% 0.08% 2.01% 0.06% 4.36 %
Infeasibility rate (1/1000) 30 0 31.7 1.11 76 5.62 110.3 22.9
Reduction in computation time 57.06% 56.61% 69.82% 63.99% 71.14% 36.56% 78.05% 55.92 %
Table 2: Simulation results for the reinforcement learning (RL) and supervised learning (SL) approaches for solving the mixed-integer linear program (5). The optimality gap is the relative distance between the objectives of the linear program (6) using the learning approaches and the objective of the mixed-integer program (5) solved to optimality with branch-and-bound. The infeasibility rate is the frequency at which the learning approaches propose discrete actions that cause linear program (6) to be infeasible. The reduction in computation time is relative to the solution time of the MILP (5).
Figure 5: Optimality gap for different prediction horizons for the reinforcement learning (RL) and supervised learning (SL) approaches in the case study.
Figure 6: Infeasibility rate for different prediction horizons for the reinforcement learning (RL) and supervised learning (SL) approaches in the case study.
Figure 7: Reduction in computation time for different prediction horizons for the reinforcement learning (RL) and supervised learning (SL) approaches in the case study.
In our case study, the dispatchable units are assumed to be identical and, consequently, a single production price is used for all of them. As a result, for optimization purposes, several discrete actions are equivalent. For instance, there is no difference in having only the first or only the third generator turned on. This allows the action space to be reduced from $322^{5}=$ 32 actions to $16$ actions, making the action space easier to explore. For the RL-based approach, Boltzmann exploration is used during training so that exploration occurs according to the decoupled Q-values. In this type of exploration, the probability of a given sub-action being chosen is proportional to the softmax of its decoupled Q-value. The policy used for exploration is described below
$\pi^{\mathrm{RL}}(\varepsilon_{\mathrm{d},l,k}(k)\ |\ \chi(k))\propto e^{\xi\cdot{Q_{l}(\chi(k),\ \varepsilon_{\mathrm{d},l,k}(k))}}$
where $\pi^{\mathrm{RL}}(\varepsilon_{\mathrm{d},l,k}(k)\ |\ \chi(k))$ represents the probability of selection of the sub-action $\varepsilon_{\mathrm{d},l,k}(k)$ in state $\chi(k)$, and $\xi$ is the exploration temperature. At the beginning of training, the temperature is chosen to $\xi=0$ so that the probabilities of selecting the sub-actions are equal, which is equivalent to random exploration. As training progresses, the temperature is gradually raised to incentivize greedier behavior.
For the SL and RL approaches, the neural network architecture consists of a single layer of an LSTM network, whose outputs are connected to a fully connected layer. The approaches are then trained with several sizes of the hidden state of the LSTM network in the set $\{64,128,256,512\}$. Next, we report the best-performing trained neural networks.
5.2 Simulation results
The simulation results for both the methods and different prediction horizons are shown in Table 2 and Figures 5, 6, and 7. The methods are compared using three performance indices namely optimality gap, infeasibility rate, and computation time. The optimality gap is the relative distance to the optimal solution computed by branch-and-bound, i.e.
$\mathrm{optimality\ gap}=\frac{J_{\mathrm{learning}}-J_{\mathrm{optimal}}}{J_{\mathrm{optimal}}}\times 100,$
where $J_{\mathrm{learning}}$ is the cost of the linear program (6) when a learning method – either RL or SL – is used to determine the discrete variables and $J_{\mathrm{optimal}}$ is the cost of the mixed-integer program (5) when it is solved to optimality with branch-and-bound. The infeasibility rate is an empirical probability of the learning methods providing discrete sequences $\epsilon_{\mathrm{d}}(k)$ that result in an infeasible linear program (6). Computation time is also relative to the time necessary for the branch-and-bound solver to reach optimality, i.e.
$\mathrm{reduction\ in\ computation\ time}=\left(1-\frac{T_{\mathrm{learning}}}{T_{\mathrm{optimal}}}\right)\times 100,$
where $T_{\mathrm{learning}}$ is the inference time of the LSTM added to the solution time of the linear program (6) and $T_{\mathrm{optimal}}$ is the solution time of the mixed-integer program (5). It is important to note that $T_{\mathrm{learning}}$ and $T_{\mathrm{optimal}}$ exclude the model-building time, which is required by the solver, as this would skew the comparison.
The simulation experiments shown in Table 2 and Figure 7 display the ability of the learning approaches to significantly reduce the online computation time of the MPC approach. The variability in the reduction in computation time stems from hyperparameter optimization (grid-search) with regard to the size of the LSTM hidden state. The computational cost for inference of the neural network decreases for smaller sizes of the hidden state, so the reported results reflect the most compact and best-performing policies. In the experiments, the SL approach performs better with a relatively small LSTM network (hidden state size is equal to $128$), so the reduction in computation is consistently lower than that of the RL approach.
The reduction in computation time comes at the price of a small optimality gap and some instances of infeasibility. Regarding feasibility, the proposed approach consistently has significantly lower infeasibility rates than that of the supervised learning method across all the considered prediction horizons. However, the supervised learning method has a considerably lower optimality gap than that of the proposed approach. This reveals a trade-off between both learning approaches in our case study: supervised learning encourages optimality and reinforcement learning promotes feasibility. Hence, the results suggest that the selection among these learning approaches ultimately depends on the relative importance of optimality and feasibility for a given application.
Concerning the scalability for an increasing prediction horizon, as seen in Figure 5, the performance of the SL method scales well for the optimality gap. Alternatively, Figure 6 reveals that the RL approach scales much better than the SL method regarding the infeasibility rate. At last, as shown in Figure 7, the reduction in computation time slightly increases for the SL method and remains at around the same level for the RL approach.
It is important to stress that our approach entirely avoids the solution of MILPs as training is only based on the solution of LPs and the MPC role lies solely in the optimization of continuous variables. Despite this apparent benefit, the training time of the RL approach was considerably longer than that of the SL approach. The structure of the underlying MILP is a key factor in determining the training time of SL since this is the main bottleneck in the training of the SL approach. Hence, the comparison between the computation time of SL and RL approaches is problem-specific.
5.3 Discussion
The results in the previous section show high potential for employing the proposed framework to efficiently control MLD systems, where a good balance between computational time, optimality, and feasibility is achieved. What follows are several points of discussion highlighting important features of the framework:
  1. 1.
    Even though discrete solutions $\epsilon_{\mathrm{d}}(k)$ that lead to the infeasibility of the LP (6) are discouraged, i.e. this event has a negative reward during training, the agent still may output an infeasible action. Moreover, feasibility can be restored either by an auxiliary control law, e.g. locally producing as much energy as possible and selling/buying the necessary power from the main grid, or by requesting a different action from the agent, e.g. selecting the second-highest Q-value instead of the highest. In the worst case, the mixed-integer linear program (5) can be solved to optimality if there is sufficient time.
  2. 2.
    In an MDP setting, some actions may not be allowed or may lead to outcomes that are not possible or feasible, depending on the state. In several implementations of RL algorithms, it is common to rule out these actions during training and inference, i.e. checking for feasibility and removing illegal actions from the action space. This typically speeds up training since the updates of the Q-function are focused only on the feasible actions. During inference, checking for feasibility can also have a positive effect in lowering the infeasibility rate. In the system considered herein, the feasibility of the action $\epsilon_{\mathrm{d}}(k)$ of the optimization problem (5) is only known after the action is sent to the environment and the LP (6) is solved. Since most of the constraints of the microgrid system depend on the continuous decision variables $\epsilon_{\mathrm{c}}(k)$, which are determined after the LP solution, it is difficult to eliminate infeasible actions before they are sent to the environment. However, even in this case, some actions can preemptively be ruled out. For instance, suppose that for a given load, the microgrid operator must turn on a minimum number of generators, and then some discrete actions that are evidently infeasible can be preemptively ruled out. Thus, implementing such a feasibility check may facilitate learning and improve performance.
  3. 3.
    It is important to stress that our approach is agnostic to the reinforcement learning algorithm that is employed. As the RL algorithm, we chose Deep Q-Learning for its simplicity. Nonetheless, numerous algorithms, e.g. Rainbow DQN, PPO, and A3C, have shown better performance than Deep Q-Learning in several standard RL benchmarks. Thus, there is a possibility for improvements in optimality and feasibility rate if more sophisticated RL algorithms are implemented.
  4. 4.
    In the context of the proposed control framework, a fundamental difference between reinforcement learning and supervised learning lies in their goals. Reinforcement learning aims to maximize the long-term reward, which is related to minimizing both the operation cost and infeasibility rate. Supervised learning targets at solving a classification problem, i.e. at finding a mapping from states to optimal solutions. Intrinsically, reinforcement learning explores and learns about infeasible actions, while supervised learning narrowly focuses on predicting the optimal action. As a result, the output of reinforcement learning is richer in providing more information about the actions and their feasibilities. On the other hand, supervised learning is much more effective at its intended purpose of predicting the optimal solution at the cost of a higher infeasibility rate.
  5. 5.
    The microgrid model used in this paper has several simplifications when compared to a more realistic one as described in [1]. The implementation of the RL-based method proposed herein in more complex systems might reveal more about its advantages and drawbacks. Moreover, due to the similarity of the control problems, our approach has the potential to be applied to different energy systems and to other meaningful infrastructure systems such as transportation systems.
  6. 6.
    Although the proposed approach was used for the economic dispatch problem of microgrids, we highlight that it can be applied to any problem of the form (5), i.e. parametric mixed-integer quadratic (linear) problems that need to be solved repeatedly. Furthermore, our approach is easily extendable to other classes of problems, such as parametric mixed-integer nonlinear programs.
6 Conclusions and future work
This paper has presented a novel approach that combines reinforcement learning and model predictive control for mixed-logical dynamical systems. The main motivation for embedding learning into the optimization of the MPC controller is to reduce its online computational complexity. The core of our approach is to decouple the decisions on the discrete and continuous optimization variables, where the RL agent is responsible for computing the discrete variables and MPC optimizes over the continuous variables. Once the discrete variables are determined by RL, the MPC control problem is simplified from a mixed-integer quadratic (linear) program into a quadratic (linear) program, which significantly improves the required solution time. The approach has been implemented to solve the economic dispatch problem for a microgrid system and it has been compared against a branch-and-bound solver and a method that uses supervised learning to decouple the decision variables. Simulation results have shown that the RL agent, whose training is based on decoupled Q-functions, outputs policies that are close to optimality, showing that the decoupling does not hinder learning. Moreover, the computational time has been substantially decreased by the learning methods relative to the branch-and-bound solver. A comparison of the reinforcement learning and supervised learning techniques has revealed an interesting trade-off between the optimality gap and the infeasibility rate of the simplified linear program. While supervised learning favors performance, reinforcement learning outperforms in terms of feasibility.
The proposed approach is particularly targeted at reducing the computational burden relative to the growth of discrete decision variables caused by the increase of the prediction horizon. However, the complexity regarding the number of actions per time step is not addressed in our work. In this context, a topic of future work will be the investigation of methods to simplify the problem at each time step, to mitigate the overall combinatorial growth of the sequence of actions. Moreover, future work will also explore methods to further reduce the computation time, the optimality gap, and the infeasibility rate. Some possible directions to improve the performance have already been discussed in the previous section. The proposed framework can also be extended to control problems that require the solution of mixed-integer nonlinear programs, which paves the way for its application in even more complex systems.
CRediT authorship contribution statement
Caio Fabio Oliveira da Silva: Conceptualization, Methodology, Software, Validation, Investigation, Data Curation, Visualization, Writing – original draft, Writing – review & editing. Azita Dabiri: Conceptualization, Methodology, Supervision, Writing – review & editing. Bart De Schutter: Conceptualization, Methodology, Supervision, Writing – review & editing, Project administration, Funding acquisition.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The code and data used in this paper are publicly available at https://github.com/fabcaio/microgrid_RL.
Acknowledgments
This research has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101018826 - CLariNet).
Appendix
Appendix A System description
In the case study, we address the economic dispatch problem in a microgrid system. This appendix describes the constituent elements of the microgrid, shown in Figure 4, and their modeling as a mixed-logical dynamical (MLD) system. This modeling framework is used for its capacity to capture the behavior of continuous and discrete dynamics and decision variables. Besides, MLD modeling paves the way for the formulation of the economic dispatch problem as a mixed-integer program using an MPC approach. The use of MLD modeling and MPC for the microgrid operation optimization was explored in [1], where the problem is cast as a mixed-integer linear program. The modeling of the microgrid was further simplified in [33] and [9] for the design of a ruled-based control policy and a learning-based control rule, respectively. Herein we use the same simplified microgrid modeling framework to assess the performance of the proposed approach.
A.1 Storage unit
The energy storage unit is described by the following equations:
$x_{\mathrm{b}}(k+1)=\begin{dcases}x_{\mathrm{b}}(k)+\frac{T_{\mathrm{s}}}{\eta_{\mathrm{d}}}P_{\mathrm{b}}(k)&\mathrm{if}\ P_{\mathrm{b}}(k)<0\\ x_{\mathrm{b}}(k)+T_{\mathrm{s}}\eta_{\mathrm{c}}P_{\mathrm{b}}(k)&\mathrm{if}\ P_{\mathrm{b}}(k)\geq 0\end{dcases}$ (9)
where $x_{\mathrm{b}}(k)$ is the energy level in the storage unit at time step $k$, $\eta_{\mathrm{c}}$ and $\eta_{\mathrm{d}}$ are the charging and discharging efficiencies, $P_{\mathrm{b}}(k)$ is power exchanged with the storage unit at time step $k$ and $T_{\mathrm{s}}$ is the sampling time of the discrete-time system. At a given time step, the storage unit can be either discharging or charging, depending on the sign of $P_{\mathrm{b}}(k)$. In order to capture this hybrid behavior, we model the storage unit as a mixed-logical dynamical (MLD) system [26]. A binary variable $\delta_{\mathrm{b}}(k)$ is introduced to signal whether the storage unit is charging or discharging, $\delta_{b}(k)=1\iff P_{\mathrm{b}}(k)\geq 0$ and $\delta_{b}(k)=0\iff P_{\mathrm{b}}(k)<0$, respectively. By defining a continuous auxiliary variable $z_{\mathrm{b}}(k)=\delta_{\mathrm{b}}(k)P_{\mathrm{b}}(k)$, (9) can be simplified and conveniently written as a single linear equation:
$x_{\mathrm{b}}(k+1)=x_{\mathrm{b}}(k)+T_{\mathrm{s}}(\eta_{\mathrm{c}}-\frac{1}{\eta_{\mathrm{d}}})z_{\mathrm{b}}(k)+\frac{T_{\mathrm{s}}}{\eta_{\mathrm{d}}}P_{\mathrm{b}}(k)$
A.2 Generation units
In the microgrid, energy can be locally produced either by dispatchable units or by renewable energy sources. The dispatchable units can be turned on/off and their power output levels can be arbitrarily chosen by the microgrid operator within operating constraints. Concerning the renewable energy sources, solar and wind energy sources are considered.
The cost for locally producing energy at time step $k$ is
$C_{\mathrm{prod}}(k)=c_{\mathrm{prod}}(k)\sum_{i=1}^{N_{\mathrm{gen}}}P_{i}^{\mathrm{dis}}(k)$ (10)
where $c_{\mathrm{prod}}(k)$ is the cost of producing energy at time step $k$, $P_{i}^{\mathrm{dis}}(k)$ represents the power generated by dispatchable unit $i$ at time step $k$, and $N_{\mathrm{gen}}$ is the total number of dispatchable units. Besides, a binary variable $\delta_{i}^{\mathrm{dis}}(k)$ is introduced to represent whether the generator $i$ is turned on (1) or off (0). This will be useful in the next section when the operating constraints are introduced and the optimization problem is formulated.
Renewable energy sources are excluded from the expression (10) because we assume zero cost for utilizing them when available. Moreover, due to their nature, the power generated by renewable sources is not controllable. The total power generated by renewable sources at time step $k$ is represented by $P_{\mathrm{res}}(k)$.
A.3 Main grid
At any given time step, the microgrid can buy or sell energy from the main grid. The power exchange at time step $k$ is represented by $P_{\mathrm{grid}}$(k). If this variable is nonnegative, the microgrid is set to import energy from the main grid. If it is negative, the microgrid is in export mode. Then, from the microgrid operator’s perspective, at time step $k$, the operation cost is represented as
$C_{\mathrm{grid}}(k)=\begin{dcases}c_{\mathrm{sell}}(k)P_{\mathrm{grid}}(k)&\iff P_{\mathrm{grid}}(k)<0\\ c_{\mathrm{buy}}(k)P_{\mathrm{grid}}(k)&\iff P_{\mathrm{grid}}(k)\geq 0\end{dcases}$ (11)
where $c_{\mathrm{sell}}(k)$ and $c_{\mathrm{buy}}(k)$ are the prices for selling and buying energy to/from the main grid, respectively, at time step $k$.
Consider the discrete auxiliary variable $[\delta_{\mathrm{grid}}(k)=1]\iff[P_{\mathrm{grid}}(k)\geq 0]$ and the continuous auxiliary variable $z_{\mathrm{grid}}(k)=\delta_{\mathrm{grid}}(k)P_{\mathrm{grid}}(k)$. Now, with the use of MLD modeling, the operation cost between the microgrid and the main grid can be expressed in a single linear equation
$C_{\mathrm{grid}}(k)=c_{\mathrm{buy}}(k)z_{\mathrm{grid}}(k)-c_{\mathrm{sell}}(k)z_{\mathrm{grid}}(k)+c_{\mathrm{sell}}(k)P_{\mathrm{grid}}(k)$ (12)
Note that the operation cost is negative in export mode, i.e. the microgrid operator profits by selling energy to the grid.
A.4 Assumptions
We only consider uncontrollable loads, i.e. the microgrid operator does not affect the power demanded by them. Furthermore, at the current operation time, the actual load and its forecast over the future are assumed to be known. Similarly, the current power generated by renewable energy sources and its forecast are assumed to be known. These are not strong assumptions, since all of these values can be estimated with the use of historical data. We also assume knowledge of the market energy prices for buying and selling energy from the main grid and for locally producing energy with the dispatchable generators.
A.5 Control scheme
We consider a hierarchical control structure. A high-level controller is concerned with planning the operation schedule: i) which generation units are turned on and their power outputs, ii) whether the storage unit is charging or discharging and the corresponding power exchange, and iii) the mode of operation of the microgrid with respect to the main grid – importing or exporting – and the amount of power flowing between them. The variables determined by the high-level controller are reported in Table 1. A low-level controller is responsible for keeping voltage, frequency, and phase within the operation range and it operates in a faster time scale. In this scheme, the high-level controller is responsible for generating the set points for the microgrid elements so that the load demand is satisfied and the operation cost is minimized. On the other hand, the low-level controller has the objective of tracking these set-points computed by the high-level controller. In this work, we assume that the low-level controller is already in place and we focus exclusively on the design of the high-level controller.
A.6 Control problem
The operation cost is defined as the sum of the the costs for locally producing energy and the costs for exchanging power with the main grid over a prediction horizon $N_{\mathrm{p}}$. Accordingly, the cost is defined as follows
$\begin{split}&J(\textbf{x}_{\mathrm{b}}(k),\gamma(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k))=\\ \sum_{l=0}^{N_{\mathrm{p}}-1}&\left(C_{\mathrm{prod}}(k+l)+C_{\mathrm{grid}}(k+l)\right)=c^{\mathrm{T}}\epsilon_{c}(k)\end{split}$ (13)
where $\textbf{x}_{\mathrm{b}}(k)=[x_{\mathrm{b}}^{\mathrm{T}}(0),\ldots,x_{\mathrm{b}}^{\mathrm{T}}(N)]^{\mathrm{T}}$, $C_{\mathrm{prod}}(k)$ and $C_{\mathrm{grid}}(k)$ are defined in (11) and (10), $c$ is a weighting vector containing the market energy prices use to compactly represent the linear dependence of the operation cost on the decision variables and $\epsilon_{\mathrm{c}}(k)$ is a stacked vector with the continuous decision variables over the prediction horizon.
Note that the cost function $J(\cdot)$ is linear with respect to $\epsilon_{\mathrm{c}}(k)$ and that it can be negative in the scenario where the microgrid operator profits from selling energy to the main grid. With the objective function defined, the optimization problem for the MPC controller can now be addressed. Along with minimizing the operation cost, the microgrid operator has also to satisfy some operating constraints, which are discussed next. Consider the following finite-horizon optimal control problem relative to the MPC controller:
$\min_{\textbf{x}_{\mathrm{b}}(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k)}$ $\ J(\textbf{x}_{\mathrm{b}}(k),\gamma(k),\epsilon_{\mathrm{c}}(k),\epsilon_{\mathrm{d}}(k))$
subj. to $x_{\mathrm{b}}(k+l+1)=x_{\mathrm{b}}(k+l)+$ (14a)
$\hskip 14.22636pt+T_{\mathrm{s}}(\eta_{\mathrm{c}}-\frac{1}{\eta_{\mathrm{d}}})z_{\mathrm{b}}(k+l)+\frac{T_{\mathrm{s}}}{\eta_{\mathrm{d}}}P_{\mathrm{b}}(k+l),$
$E_{2}\delta(k+l)+E_{3}z(k+l)\leq E_{1}u(k+l)+$ (14b)
$\hskip 28.45274pt+E_{4}x_{\mathrm{b}}(k)+E_{5},$
$P_{\mathrm{b}}(k+l)=\sum_{j=1}^{N_{\mathrm{gen}}}(P_{j}^{\mathrm{dis}}(k+l))+$ (14c)
$+P_{\mathrm{res}}(k+l)+P_{\mathrm{grid}}(k+l)-P_{\mathrm{load}}(k+l),$
$\underline{P}_{\mathrm{b}}\leq P_{\mathrm{b}}(k+l)\leq\bar{P}_{\mathrm{b}},$ (14d)
$\underline{P}_{\mathrm{grid}}(k+l)\leq P_{\mathrm{grid}}(k+l)\leq\bar{P}_{\mathrm{grid}},$ (14e)
$\delta_{j}^{\mathrm{dis}}(k+l)\underline{P}_{j}^{\mathrm{dis}}\leq P_{j}^{\mathrm{dis}}(k+l)\leq\delta_{j}^{\mathrm{dis}}(k+l)\bar{P}_{j}^{\mathrm{dis}},$ (14f)
for $\text{for }j=1,\ldots,N_{\mathrm{gen}},\ \text{and}$
for $\text{for }l=0,\dots,N_{\mathrm{p}}-1$
where the cost $J(\cdot)$ is defined by (13) and the decision variables and parameters are described in Tables 1 and 3. The first constraint (14a) concerns the dynamic of the storage unit. The equations (14b) arise from the introduction of the discrete and auxiliary variables introduced for the MLD modeling. The power balance of the microgrid and its connection with the main grid is enforced by constraint (14c) – the generation must correspond to the load at all time steps. The lower and upper limits for the power exchange with the storage unit and with the main grid are represented in (14d) and (14e), respectively. At last, the dispatchable generation units have minimum and maximum operation levels which are represented by (14f).
The objective function and the constraints of (14) are linear with respect to the optimization variables. Since the model and the constraints are linear, they can be conveniently represented in MLD form as in (1). As a result, based on the general formulation (5), the problem can be recast as a mixed-integer linear problem, which is represented below
$\begin{split}V(x_{\mathrm{b}}(k),\ \gamma(k))=\min_{\epsilon_{d}(k),\ \epsilon_{c}(k)}\ &c_{c}^{\mathrm{T}}\epsilon_{c}(k)\\ \text{subj. to }&G[\epsilon_{\mathrm{c}}^{\mathrm{T}}(k),\ \epsilon_{\mathrm{d}}^{\mathrm{T}}(k)]^{\mathrm{T}}\leq w+Sx_{\mathrm{b}}(k)\end{split}$ (15)
Parameter Description Value
$\bar{x}_{\mathrm{b}}$ maximum storage unit level 250 [kWh]
$\underline{x}_{\mathrm{b}}$ minimum storage unit level 25 [kWh]
$\bar{P}_{\mathrm{grid}}$ maximum power exchange with main grid 1000 [kW]
$\underline{P}_{\mathrm{grid}}$ minimum power exchange with main grid -1000 [kW]
$\bar{P}_{\mathrm{b}}$ maximum power exchange with storage unit 100 [kW]
$\bar{P}_{\mathrm{b}}$ minimum power exchange with storage unit -100 [kW]
$\bar{P}_{i}^{\mathrm{dis}}$ maximum power output of dispatchable unit $i$ 100 [kW]
$\underline{P}_{i}^{\mathrm{dis}}$ minimum power output of dispatchable unit $i$ 100 [kW]
$\eta_{c}$ charging efficiency of storage unit 0.9
$\eta_{d}$ discharging efficiency of storage unit 0.9
$N_{gen}$ number of generators 3
Table 3: Microgrid parameters, their descriptions, and their values.
Appendix B Supervised learning approach
The training procedure of the SL approach presented in [12] is briefly summarized to highlight fundamental differences between both RL and SL approaches. The approach presented in [12] employs an LSTM network for classification, which is the same neural network architecture used in our RL approach for the policy approximation, facilitating the comparison between the performances of both methods. Let the optimal solution of the MPC problem in (15) for a given state $\chi(k)$ be defined by the tuple $(\epsilon_{d}^{*}(k),\ \epsilon_{c}^{*}(k))$ containing the discrete and continuous optimization variables. The dataset for training is created by solving the MPC problem to optimality with branch-and-bound a number $N_{\mathrm{data}}$ of times with different initial augmented states, which are randomly sampled across the system’s operating region of the state-space. The resulting pairs $\{(\chi(k),\ \epsilon_{d}^{*}(k))\}_{k=1}^{N_{\mathrm{data}}}$ are stored in a data buffer $\mathcal{D}$. Note that the continuous optimal solutions are not stored in the data buffer because the learning objective only concerns the prediction of the discrete optimal solutions. Having the dataset, it is straightforward to tune the parameters of the LSTM network by gradient descent on the following expression:
$L=\frac{1}{T}\sum_{k=1}^{T}\sum_{l=0}^{N_{\mathrm{p}}-1}g\left((y_{l}(k)-[\phi_{\theta}^{\mathrm{LSTM}}(\chi(k))]_{l})\right)$
where $g(\cdot)$ is a loss function, e.g. cross entropy function, $T$ is the size of the mini-batch, the target $y_{l}(k)$ is defined as the one-hot encoding of the optimal sub-action $\varepsilon^{*}_{\mathrm{d},l}(k)$ and the $l$-th output of the unrolled LSTM network with weights $\theta$ is denoted by $[\phi_{\theta}^{\mathrm{LSTM}}(\cdot)]_{k}$. For more details, the reader is referred to [12].